# References.bib Combining Hybrid and Opaque Scintillator Techniques in Search for Double Beta Plus Decays A Concept Study for the NuDoubt++ Experiment

### v1

@@ 1 Introduction
Double beta searches have been proposed individually for hybrid and opaque scintillation detectors a few years ago. Early ideas in 2016 involved the use of prompt directional Cherenkov light in addition to isotropic scintillation light to discriminate double electron decays from solar B-8 neutrino background [ELAGIN_2017, Jiang_2019]. In 2022, the approach was advanced to the usage of slow scintillators that improve separation of Cherenkov and scintillation light [Dunger_2022]. In the same year, our idea to exploit the ratio of Cherenkov and scintillation light in an hybrid detector for the search of double positron emission was presented [Wurm_2022]. Before, sensitivity towards in a kilotonne-scale hybrid detector was shown in a simulation study by the Theia collaboration in 2020 [Theia_2019]. First demonstrations of the hybrid detector concept through small-scale prototypes were published since 2017 [Caravaca_2017, Caravaca_2020]. A first tonne-scale hybrid detector for double electron decays was put forward by the NuDot collaboration in 2018 based on earlier small-scale prototypes [NuDot_2018]. Also since 2018, the LiquidO consortium proposed for the first time searches for double weak decays exploiting distinct event topologies in opaque media suitable for double electron and double positron searches [LiquidO_2018_beta, LiquidO_2023_beta, LiquidO_2023_beta_2, LiquidO_2023_beta_3]. This idea is refined by the LiquidO collaboration [LiquidO_2024_beta]. Demonstrations of the opaque detector concept through small-scale prototypes were published since 2021 [LiquidO_2021, LiquidO_2024_LIME, LiquidO_2024].
The present article discusses both the general properties of a new hybrid-opaque experiment and a first outline for a tonne-scale (scintillator mass) demonstrator experiment. The corresponding detector features a fiducial volume of one cubic metre of hybrid-slow opaque scintillator densely instrumented with novel optimised wavelength-shifting fibers. We show that such a detector is well suited to perform first measurements of currently little explored two-neutrino double weak decays, such as double positron emission (2) and simultaneous electron capture and positron emission (EC), and set limits on the corresponding neutrinoless EC and 2 half-lives. The isotopes in question are Kr-78, Xe-124, and Cd-106. Thus, the prototype would provide both a demonstration of the essential detector properties and additional data points to support theoretical modeling of nuclear matrix elements. If successful, and decays would be the rarest standard model (SM) decay processes ever observed.
@@ 2 Motivation of Searches for Double Beta Plus Decays
The limited exploration of these modes arises from the suppressed decay probabilities, less favourable decay Q-values, experimental challenges in the observation of a signature of these decay modes and low natural abundances of suitable candidate nuclei [Graf_2022, Blaum_2020, Suhonen_2013, Belli_2019]. The identification of 2EC can be achieved by detecting the cascade of X-rays and Auger electrons generated during the filling of vacancies subsequent to the capture of two atomic electrons. The nuclear binding energy, Q, released during this process (typically around 1 MeV) is predominantly carried away by the two neutrinos, which remain undetected within the detector. For decay modes associated with positrons, the experimental signature is the simultaneous emission of one (in case of EC) or two (in case of ) pairs of 511 keV annihilation gamma-rays.
@@ 3.3 Combining Hybrid and Opaque Scintillators
The combination of opaque and hybrid media becomes possible through the recent progress of extending NoWaSH to cover a variety of solvents, including those of the hybrid slow solvent scintillator, through the use of novel wax types [Schoppmann_2024]. Since our detector medium is hybrid opaque NoWaSH, high levels of isotope loading become possible. Usually, such high isotope loading would degrade the transparency of a classic scintillator to a point where it becomes unusable [Buck_2016]. However, in an opaque medium, the requirements on transparency are vastly relaxed due to local light read-out in combination with a scattering length of about 1 mm, which leads to much shorter average paths lengths of optical photons [Buck_2019, LiquidO_2021]. We describe our strategy for isotope loading in section 4.
@@ 3.4 Expected Performance
Despite our background discrimination strategy based on combined hybrid opaque scintillator, in the beyond standard model cases of and , the most important background rejection strategy is the definition of a very small energy window, the region of interest (ROI), around the mono-energetic line from the positron(s). To achieve a small ROI, the energy resolution of the detector, which widens the mono-energetic lines of neutrinoless decay modes to peaks, has to be very good. Based on the preparatory work and the considerations made above, we propose a hybrid-slow opaque scintillator consisting of 88 wt.% linear alkylbenzene (LAB), 10 wt.% diisopropylnaphtalene (DIN), and 2 wt.% wax with diphenyloxazole (PPO) at a concentration of 1.0 g/l as wavelength shifter. This scintillator achieves a light yield of 9000 photons/MeV, an attenuation length of 17 m, and a scattering length of 2 mm [Buck_2019, Schoppmann_2024, Steiger_2024_slow, Stock_2024]. Due to its slow scintillation emission with primary (secondary) decay-time constant of 13 ns (26 ns), the scintillation light is strongly delayed with respect to the Cherenkov peak. The scintillator achieves a purity of the Cherenkov light in the Cherenkov peak of 80% [Steiger_2024_slow, Stock_2024]. Current opaque detector readout technology is expected to reach more than 500 PE/MeV (Photon Equivalences) [LiquidO_2024]. The largest loss of light happens at the interfaces between scintillator and wavelength-shifting fibres, as the geometrical capture efficiency via total internal reflection of the isotropically emitted shifted light is under 10%. We plan to improve this loss by using optimised wavelength-shifting light guides (OWL-fibres). These OWL-fibres feature an up to four times increase capture efficiency due to an optimised geometrical distribution of wavelength-shifter in the light guide [JOHNSON_1994] and are detailed in section 5. Thus, a light yield of more than 800 PE/MeV appears possible in NuDoubt++.
@@ 4 Isotope Candidates
In the search for beta plus decays, three isotopes are of particular interest: Kr-78, Cd-106, and Xe-124 (cf. Table 1). All have relatively high Q-values well above the gamma-line from Tl-208 at 2.614 MeV leading to low background rates. As for the half-lives of these isotopes concerning neutrinoless and two-neutrino decay channels involving positron emission, only lower limits of to years have been established thus far. In the following, we discuss loading for the noble gases krypton and xenon in subsection 4.1 and for the metal cadmium in subsection 4.2.
@@ 4.1 Noble Gases
Among the noble gases, krypton and xenon stand out as promising candidates for double beta decay studies. Krypton features two double beta isotopes and has been explored much less than xenon in this context. Kr-84 is a isotope with low Q-value and is henceforth neglected in this article. Kr-78 decays to Se-78 via double weak interaction with a Q-value of 2.881 MeV. The expected half-lives for the two-neutrino 2EC and EC decays in this isotope are about years, making the latter readily detectable in a small NuDoubt++ detector. Two-neutrino has an expected half-live of years, being reachable in an upgraded version of NuDoubt++. The expected half-lives for the neutrinoless double beta decay modes start from about years for effective neutrino masses of few tenths of eV [Suhonen_2013].
Similarly, xenon has two double beta isotopes. Xe-136 is a isotope with a Q-value of 2.459 MeV and heavily explored by existing experiments [Gando_2012, Xenon_2022]. Xe-124 decays to Te-124 via double weak interaction with a slightly lower Q-value of 2.857 MeV. The mode has been recently observed with a half-live of years [Aprile_2019]. The expected half-lives for the two-neutrino 2EC and EC decays are in the range of years, indicating that EC could also be detected in our experiment. The expected half-lives for the neutrinoless double beta decay modes start from years [Suhonen_2013_Xe].
@@ 4.2 Cadmium
The isotope Cd-106 is another suitable candidate for the search for positive double weak inteactions. It decays into Pd-106 and has a high Q-value of 2.775 MeV. The two-neutrino 2EC and EC double beta decays have expected half-lives of years and is thus quickly detectable in our proposed experiment. Two-neutrino double beta decay has an expected half-life of [SUHONEN_2001]. The half-lives for the neutrinoless double beta decay modes start from years [SUHONEN_2011].
@@ 6.4 Backgrounds from Detector Structures and Laboratory
External gamma rays from the surroundings of the detector and its outer layers impose a further background for rare-decay experiments like NuDoubt++ . Especially the high-energy gammas of Tl-208 and Bi-214 (daughter nuclei of U-238 and Th-232) can generate signal-like events around the Q-value. During the detector design phase, material selection necessitates thorough radiopurity characterisation, with particular attention to materials in direct contact with the source, such as the vessels.
@@ 6.5 Background Discrimination
Figure 5 illustrates the predicted event rates and spectra for relevant liquid scintillator backgrounds and the two-neutrino signals of krypton, assuming a detector placement at the Laboratori Nazionali del Gran Sasso (LNGS) (overburden of 3800 metre water equivalent [Miramonti_2005]). In this spectrum, event discrimination between electrons and gammas on the one hand and positron events on the other hand have been taken into account by applying reduction factors. Opaque detector technology alone is expected to achieve background suppression factors of up to based on simple counting of blobs [LiquidO_2021]. Through combination of opaque and hybrid detection techniques, we expect to significantly improve the suppression factors based on the clear separation of event categories via their Č/S ratio [Mossel_2024] as shown in Figure 6 and based on the additional novel discrimination strategies discussed in section 3.
@@
In the optimistic scenario, we study cut efficiencies for opaque and hybrid techniques based on Monte-Carlo simulations. Here, we apply a clustering algorithm derived from the Cambridge-Aachen jet clustering algorithm [Bentvelsen_1998] to identify blobs and discriminate events based on their number of blobs. We further consider the Č/S ratio of event types, taking into account the response of your slow scintillator and fibres in a detector simulation [Mossel_2024].
@@ 7 Design of Detector Prototype
Based on the concepts discussed above, we consider here a prototype detector of about one metric tonne hybrid-slow opaque scintillator. This demonstrator can first be used to verify the opaque and hybrid particle discrimination capabilities of the new concept, as well as the complementarity of both discrimination strategies. For this purpose, the prototype will offer the possibility to deploy positron, gamma, and electron sources in its centre. We foresee a detector with cylindrical shape of equal diameter and height of roughly 110 cm. OWL-fibres will be running parallel to its symmetry axis, as illustrated in Figure 7. While a spherical detector geometry offers the best volume to surface ratio, we propose a cylindrical geometry to ease the installation of OWL-fibres. Each OWL-fibre is connected with a SiPM on both ends to allow for a high light collection efficiency and therefore energy resolution, as well as positional resolution along the OWL-fibre exploiting signal delays between both ends. The arrangement of OWL-fibres just along one direction is sufficient for effective particle discrimination due to the complexity of the topological pattern of positrons compared to gammas or electrons, as shown in Figure 2. The OWL-fibres are arranged in a triangular grid with a pitch of between nearest neighbours, resulting in about 13,000 fibres in total. At this spacing, an average blob will illuminate about 15 OWL-fibres and yield an energy resolution of 6%. From Monte Carlo simulations, we see that a cylindrical detector of one metric tonne is able to fully contain all electron events, but only 50% of positron and gamma events starting at its centre. Nevertheless, it is possible to reconstruct the full event energy of all positrons and gammas even in this case because the original energy of an annihilation gamma can be determined just from the angle of the first Compton-scatter and the energy deposited in the first Compton-scatter through Equation 6, as explained in subsection 3.2. The first two Compton-scatter vertices of annihilation gammas are contained for 97% of positrons in our detector, as we see in Monte Carlo simulations. In addition, the kinetic energy of positrons and electrons is always deposited in a single blob at the location of their creation inside the detector. The detector could be surrounded by an active veto with good particle ID, especially to tag neutrons [DoubleChooz_2022_detector, AMOTech_2023, Buck_2019_Stereo].
@@ 8 Sensitivity
Similar to the sensitivity study for the two-neutrino decay, a lower limit can be derived for the neutrinoless case. To do this, we are only considering signal events at the Q-value of this decay at 2.881 MeV and compare them to the background in the ROI given by our energy resolution. For this second study, we use the optimistic discrimination factors from Table 3 and the optimistic background rates for our OWL-fibres from Table 2. As seen in Figure 9, with a detector loaded with 50% enriched krypton at 5 bar overpressure in a 10 kg (scintillator mass) central fiducial vessel, we can set a limit on on the order of years at 90% C.L. after an exposure of one tonne-week or 20 kg-years. We can achieve a similar limit for the decay mode. Our background index inside the ROI lies at . We can further improve the background index to by applying stricter cuts at the expense of overall signal rate.
@@ Acknowledgements
Following the topical order of the article, we thank Lukáš Gráf for discussions on phenomenological aspects of double beta plus searches. We especially thank Anatael Cabrera Serra for discussion on initial ideas regarding double beta plus measurements with opaque detectors. We further thank Hardy Simgen and Jonas Westermann for discussions and support regarding low-radioactivity proportional counters, Kai Loo and Arshak Jafar for providing simulations used for early studies on the Č/S ratio, as well as the authors of reference [Steiger_2024_slow] (H. Steiger et al.) for granting us early access. We thank members of the IceCube, LiquidO, and Theia collaborations for feedback and discussion on various aspects of our concept.

### v2

# References.bib Combining Hybrid and Opaque Scintillator Techniques in the Search for Double Beta Plus Decays A Concept Study for the NuDoubt++ Experiment

@@ 1 Introduction
Double beta searches have been proposed individually for hybrid and opaque scintillators a few years ago, while the idea to exploit the unique signature of four or two annihilation gamma-rays for background suppression in search of double positron emission (2) or simultaneous electron capture and positron emission (EC), respectively, dates back several decades [Zeldovich_1981]. Early ideas on hybrid detection published in 2016 involved the use of prompt directional Cherenkov light in addition to isotropic scintillation light to discriminate double electron decays from solar B-8 neutrino background [ELAGIN_2017, Jiang_2019]. In 2022, an approach using slow scintillators that improve separation of Cherenkov and scintillation light was published [Dunger_2022]. In the same year, our idea to exploit the ratio of Cherenkov and scintillation light in an hybrid detector for the search of double positron emission was disseminated [Wurm_2022]. Before, sensitivity towards in a kilotonne-scale hybrid detector was shown in a simulation study published by the Theia collaboration in 2020 [Theia_2019]. First demonstrations of the hybrid detector concept through small-scale prototypes were published since 2017 [Caravaca_2017, Caravaca_2020]. A first tonne-scale hybrid detector for double electron decays was put forward by the NuDot collaboration in 2018 based on earlier small-scale prototypes [NuDot_2018]. Also since 2018, the LiquidO collaboration disseminated for the first time the concept of searches for double weak decays exploiting distinct event topologies in opaque media suitable for double electron and double positron searches [LiquidO_2023_beta_1, LiquidO_2023_beta_2, LiquidO_2023_beta_3]. This idea is refined by the LiquidO collaboration [LiquidO_2024_beta]. Demonstrations of the opaque detector concept through small-scale prototypes were published by the LiquidO collaboration since 2021 [LiquidO_2021, LiquidO_2024_LIME, LiquidO_2024].
The present article discusses both the general properties of a new hybrid-opaque experiment and a first outline for a tonne-scale (scintillator mass) demonstrator experiment. The corresponding detector features a fiducial volume of one cubic metre of hybrid-slow opaque scintillator densely instrumented with novel optimised wavelength-shifting fibers. We show that such a detector is well suited to perform first measurements of currently little explored two-neutrino double weak decays, such as 2 and EC, and set limits on the corresponding neutrinoless EC and 2 half-lives. The isotopes in question are Kr-78, Xe-124, and Cd-106. Thus, the prototype would provide both a demonstration of the essential detector properties and additional data points to support theoretical modeling of nuclear matrix elements. If successful, and decays would be the rarest standard model (SM) decay processes ever observed.
@@ 2 Motivation of Searches for Double Beta Plus Decays
The limited exploration of these modes arises from the suppressed decay probabilities, less favourable decay Q-values, experimental challenges in the observation of a signature of these decay modes and low natural abundances of suitable candidate nuclei [Graf_2022, Blaum_2020, Suhonen_2013, Belli_2019, Hirsch_1994]. The identification of 2EC can be achieved by detecting the cascade of X-rays and Auger electrons generated during the filling of vacancies subsequent to the capture of two atomic electrons. The nuclear binding energy, Q, released during this process (typically around 1 MeV) is predominantly carried away by the two neutrinos, which remain undetected within the detector. For decay modes associated with positrons, the experimental signature is the simultaneous emission of one (in case of EC) or two (in case of ) pairs of 511 keV annihilation gamma-rays.
@@ 3.3 Combining Hybrid and Opaque Scintillators
The combination of hybrid and opaque media becomes possible through the recent progress of extending NoWaSH to cover a variety of solvents, including those of the hybrid slow solvent scintillator, through the use of novel wax types [Schoppmann_2024]. Since our detector medium is hybrid opaque NoWaSH, high levels of isotope loading become possible. Usually, such high isotope loading would degrade the transparency of a classic scintillator to a point where it becomes unusable [Buck_2016]. However, in an opaque medium, the requirements on transparency are vastly relaxed due to local light read-out in combination with a scattering length of about 1 mm, which leads to much shorter average paths lengths of optical photons [Buck_2019, LiquidO_2021]. We describe our strategy for isotope loading in section 4.
@@ 3.4 Expected Performance
Despite our background discrimination strategy based on combined hybrid opaque scintillator, in the beyond standard model cases of and , the most important background rejection strategy is the definition of a very small energy window, the region of interest (ROI), around the mono-energetic line from the positron(s). To achieve a small ROI, the energy resolution of the detector, which widens the mono-energetic lines of neutrinoless decay modes to peaks, has to be very good. Based on the preparatory work and the considerations made above, we propose a hybrid-slow opaque scintillator consisting of 88 wt.% linear alkylbenzene (LAB), 10 wt.% diisopropylnaphtalene (DIN), and 2 wt.% wax with diphenyloxazole (PPO) at a concentration of 1.0 g/l as wavelength shifter. This scintillator achieves a light yield of 9000 photons/MeV, an attenuation length of 17 m, and a scattering length of 2 mm [Buck_2019, Schoppmann_2024, Steiger_2024_slow, Stock_2024]. Due to its slow scintillation emission with primary (secondary) decay-time constant of 13 ns (26 ns), the scintillation light is strongly delayed with respect to the Cherenkov peak. The scintillator achieves a purity of the Cherenkov light in the Cherenkov peak of 80% [Steiger_2024_slow, Stock_2024]. Current opaque detector readout technology is expected to reach between 350 and 550 PE/MeV (Photon Equivalences) [LiquidO_2024]. The largest loss of light happens at the interfaces between scintillator and wavelength-shifting fibres, as the geometrical capture efficiency via total internal reflection of the isotropically emitted shifted light is under 10%. We plan to improve this loss by using optimised wavelength-shifting light guides (OWL-fibres). These OWL-fibres feature an up to four times increase capture efficiency due to an optimised geometrical distribution of wavelength-shifter in the light guide [JOHNSON_1994] and are detailed in section 5. Thus, a light yield of more than 800 PE/MeV appears possible in NuDoubt++.
@@ 4 Isotope Candidates
In the search for beta plus decays, three isotopes are of particular interest: Kr-78, Cd-106, and Xe-124 (cf. Table 1). All have relatively high maximal detectable energies
in NuDoubt++, where () denotes the atomic Q-value of the () decay, the rest mass of an electron/positron, the mass difference between the neutral parent and daughter atoms, and where the atomic binding energy of electrons has been neglected. is well above the gamma-line from Tl-208 at 2.614 MeV for all isotopes, leading to low background rates. As for the half-lives of these isotopes concerning neutrinoless and two-neutrino decay channels involving positron emission, only lower limits of to years have been established thus far. In the following, we discuss loading for the noble gases krypton and xenon in subsection 4.1 and for the metal cadmium in subsection 4.2.
@@ 4.1 Noble Gases
Among the noble gases, krypton and xenon stand out as promising candidates for double beta decay studies. Krypton features two double beta isotopes and has been explored much less than xenon in this context. Kr-84 is a isotope with low Q-value and is henceforth neglected in this article. Kr-78 decays to Se-78 via positive double weak interaction with a of 2.881 MeV. The expected half-lives for the two-neutrino 2EC and EC decays in this isotope are about years, making the latter readily detectable in a small NuDoubt++ detector. Two-neutrino has an expected half-live of years, being reachable in an upgraded version of NuDoubt++. The expected half-lives for the neutrinoless double beta decay modes start from about years for effective neutrino masses of few tenths of eV [Suhonen_2013, Hirsch_1994].
Similarly, xenon has two double beta isotopes. Xe-136 is a isotope with a Q-value of 2.459 MeV and heavily explored by existing experiments [Gando_2012, Xenon_2022]. Xe-124 decays to Te-124 via positive double weak interaction with a slightly lower of 2.857 MeV. The mode has been recently observed with a half-live of years [Aprile_2019]. The expected half-lives for the two-neutrino 2EC and EC decays are in the range of years, indicating that EC could also be detected in our experiment. Two-neutrino has an expected half-live of years, being again reachable in an upgraded version of NuDoubt++ [Hirsch_1994]. The expected half-lives for the neutrinoless double beta decay modes start from years for effective neutrino masses of few tenths of eV [Suhonen_2013_Xe, Hirsch_1994].
@@ 4.2 Cadmium
The isotope Cd-106 is another suitable candidate for the search for positive double weak interactions. It decays into Pd-106 and has a high of 2.775 MeV. The two-neutrino 2EC and EC double beta decays have expected half-lives of years and is thus quickly detectable in our proposed experiment. Two-neutrino decay has an expected half-life of [SUHONEN_2001, Hirsch_1994]. The half-lives for the neutrinoless double beta decay modes start from years [SUHONEN_2011, Hirsch_1994].
@@ 6.4 Backgrounds from Detector Structures and Laboratory
External gamma-rays from the surroundings of the detector and its outer layers impose a further background for rare-decay experiments like NuDoubt++. Especially the high-energy gammas of Tl-208 and Bi-214 (daughter nuclei of U-238 and Th-232) can generate signal-like events around . During the detector design phase, material selection necessitates thorough radiopurity characterisation, with particular attention to materials in direct contact with the source, such as the vessels.
@@ 6.5 Background Discrimination
Figure 5 illustrates the predicted event rates and spectra for relevant liquid scintillator backgrounds and the two-neutrino signals of krypton, assuming a detector placement at the Laboratori Nazionali del Gran Sasso (LNGS) (overburden of 3800 metre water equivalent [Miramonti_2005]). In this spectrum, event discrimination between electrons and gammas on the one hand and positron events on the other hand have been taken into account by applying reduction factors. Opaque detector technology alone is expected to achieve background suppression factors of up to based on simple counting of blobs [LiquidO_2021]. Through combination of hybrid and opaque detection techniques, we expect to significantly improve the suppression factors based on the clear separation of event categories via their Č/S ratio [Mossel_2024] as shown in Figure 6 and based on the additional novel discrimination strategies discussed in section 3.
@@
In the optimistic scenario, we study cut efficiencies for hybrid and opaque techniques based on Monte-Carlo simulations. Here, we apply a clustering algorithm derived from the Cambridge-Aachen jet clustering algorithm [Bentvelsen_1998] to identify blobs and discriminate events based on their number of blobs. We further consider the Č/S ratio of event types, taking into account the response of your slow scintillator and fibres in a detector simulation [Mossel_2024].
@@ 7 Design of Detector Prototype
Based on the concepts discussed above, we consider here a prototype detector of about one metric tonne hybrid-slow opaque scintillator. This demonstrator can first be used to verify the hybrid and opaque particle discrimination capabilities of the new concept, as well as the complementarity of both discrimination strategies. For this purpose, the prototype will offer the possibility to deploy positron, gamma, and electron sources in its centre. We foresee a detector with cylindrical shape of equal diameter and height of roughly 110 cm. OWL-fibres will be running parallel to its symmetry axis, as illustrated in Figure 7. While a spherical detector geometry offers the best volume to surface ratio, we propose a cylindrical geometry to ease the installation of OWL-fibres. Each OWL-fibre is connected with a SiPM on both ends to allow for a high light collection efficiency and therefore energy resolution, as well as positional resolution along the OWL-fibre exploiting signal delays between both ends. The arrangement of OWL-fibres just along one direction is sufficient for effective particle discrimination due to the complexity of the topological pattern of positrons compared to gammas or electrons, as shown in Figure 2. The OWL-fibres are arranged in a triangular grid with a pitch of between nearest neighbours, resulting in about 13,000 fibres in total. At this spacing, an average blob will illuminate about 15 OWL-fibres and yield an energy resolution per blob of 6%. From Monte Carlo simulations, we see that a cylindrical detector of one metric tonne is able to fully contain all electron events, but only 50% of positron and gamma events starting at its centre. Nevertheless, it is possible to reconstruct the full event energy of all positrons and gammas even in this case because the original energy of an annihilation gamma can be determined just from the angle of the first Compton-scatter and the energy deposited in the first Compton-scatter through Equation 6, as explained in subsection 3.2. The first two Compton-scatter vertices of annihilation gammas are contained for 97% of positrons in our detector, as we see in Monte Carlo simulations. In addition, the kinetic energy of positrons and electrons is always deposited in a single blob at the location of their creation inside the detector. The detector could be surrounded by an active veto with good particle ID, especially to tag external backgrounds [DoubleChooz_2022_detector, AMOTech_2023, Buck_2019_Stereo].
@@ 8 Sensitivity
Similar to the sensitivity study for the two-neutrino decay, a lower limit can be derived for the neutrinoless case. To do this, we only consider signal events around 2.881 MeV and compare them to the background in the ROI. We define the ROI between 2.82 and 3.00 MeV. For this second study, we use the optimistic discrimination factors from Table 3 and the optimistic background rates for our OWL-fibres from Table 2. As seen in Figure 9, with a detector loaded with 50% enriched krypton at 5 bar overpressure in a 10 kg (scintillator mass) central fiducial vessel, we can set a limit on on the order of years at 90% C.L. after an exposure of one tonne-week or 20 kg-years. Our background index inside the ROI, not considering signal efficiencies, lies at . We can further improve the background index to by applying stricter cuts at the expense of overall signal rate. We can achieve a similar limit for the decay mode.
@@ Acknowledgements
Following the topical order of the article, we thank Lukáš Gráf for discussions on phenomenological aspects of double beta plus searches. We especially thank Anatael Cabrera for discussion on initial ideas regarding double beta plus measurements with the LiquidO technology. We further thank Hardy Simgen and Jonas Westermann for discussions and support regarding low-radioactivity proportional counters, Kai Loo and Arshak Jafar for providing simulations used for early studies on the Č/S ratio, as well as the authors of reference [Steiger_2024_slow] (H. Steiger et al.) for granting us early access. We thank members of the IceCube, LiquidO, and Theia collaborations for feedback and discussion on various aspects of our concept.